The length of a rectangle is twice the width of the rectangle. The length of a diagonal of the rectangle is 25cm. How do I find out the are of the rectangle?
Answer by Greenfinch(383)

If the width is X, the length is 2X and the diagonal is sqrt ((X^2) + (2X)^2)
So sqrt(5X^2) = 25 or 5X^2 = 625
X^2 = 125
Area is X x 2X = 2X^2 = 250